# Joint User Association and Resource Allocation for Multi-Cell Networks with Adaptive Semantic Communication

### 05 Jan 2024 v2

@@ Abstract
Semantic communication is a promising communication paradigm that utilizes Deep Neural Networks (DNNs) to extract the information relevant to downstream tasks, hence significantly reducing the amount of transmitted data. In current practice, the semantic communication transmitter for a specific task is typically pre-trained and shared by all users. However, due to user heterogeneity, it is desirable to use different transmitters according to the available computational and communication resources of users. In this paper, we first show that it is possible to dynamically adjust the computational and communication overhead of DNN-based transmitters, thereby achieving adaptive semantic communication. After that, we investigate the user association and resource allocation problem in a multi-cell network where users are equipped with adaptive semantic communication transmitters. To solve this problem, we decompose it into three subproblems involving the scheduling of each user, the resource allocation of each base station (BS), and the user association between users and BSs. Then we solve each problem progressively based on the solution of the previous subproblem. The final algorithm can obtain near-optimal solutions in polynomial time. Numerical results show that our algorithm outperforms benchmarks under various situations.
Index Terms:
Semantic communication, resource allocation, user association, multi-cell networks
@@ I Introduction
It has been over seventy years since the inception of Shannon’s information theory [1]. During this period, traditional wireless communication technologies, aiming at the accurate transmission of bits, have made significant advancements and approached the Shannon limit. However, with the emergence of modern applications such as virtual reality and smart cities, the amount of data that needs to be transmitted continues to grow at an unprecedented pace. To meet the transmission demands with limited communication resources, a new communication paradigm called semantic communication has been proposed [2, 3, 4]. With the help of artificial intelligence technologies, semantic communication only transmits the information relevant to downstream tasks (which is usually referred to as semantic information), hence leading to a significant reduction in data traffic.
Although semantic communication can reduce the amount of transmitted data, extracting semantic information introduces additional computational overhead. In current practice, users performing the same task typically use the same semantic communication transmitter. However, due to the heterogeneity of users and networks, it is desirable to dynamically adjust the transmitter according to the available computational and communication resources. For example, since the transmitters are implemented by Deep Neural Networks (DNNs), we can decrease the model size of DNNs for users with limited computational capabilities. Similarly, for users with poor channel conditions, we can further reduce the amount of transmitted data by only transmitting a portion of the semantic information. By combining these two methods, we can obtain a range of semantic communication schemes with different computational overhead, communication overhead, and result accuracies. By selecting the optimal semantic communication scheme for each user, we can achieve adaptive semantic communication and significantly improve the system performance.
In this paper, we consider the joint user association and resource allocation problem for multi-cell networks with adaptive semantic communication. However, due to user heterogeneity (e.g. variations in computational capabilities, maximum transmitting power, and energy budget) and network heterogeneity (e.g. variations in user distribution, base station coverage, and channel conditions), there does not exist a straightforward solution for this problem. To solve this problem, we need to address the following challenges.
  • Firstly, both computation and communication processes consume energy and time. Therefore, there exists a trade-off between the computational workload and the communication workload that a user can handle. Since the result accuracy depends on both kinds of workloads, we need to quantify this trade-off in order to obtain the optimal semantic communication scheme. However, how to quantify this trade-off is not explicitly clear.
  • Secondly, in addition to transmitting power and transmission time, the spectrum bandwidth also influences the amount of transmitted data, thereby further affecting the result accuracy. However, due to user heterogeneity, the impact of spectrum bandwidth on result accuracy varies from user to user and hard to quantify. As a result, there is no simple rule to find the optimal spectrum allocation.
  • Thirdly, due to network heterogeneity, associating all users with the nearest base station (BS) may result in severe load imbalance. Therefore, to make appropriate user association decisions, we need to consider both load balancing and channel conditions comprehensively. However, it is not clear how to measure the impact of these two factors.
To address the above challenges, this paper designs an efficient approximate algorithm to obtain near-optimal user association and resource allocation decisions in polynomial time. Our main contributions can be summarized as follows:
  • We propose the concept of adaptive semantic communication and formulate the user association and resource allocation problem in a multi-cell network. We also show the formulated problem is strongly NP-hard.
  • We progressively solve the problem by decomposing it into three subproblems involving the scheduling of each user, the resource allocation of each BS, and the user association between users and BSs. The solution to each subproblem relies on the properties of the optimal solution of the previous subproblem. The final algorithm can obtain near-optimal solutions in polynomial time.
  • We conduct extensive simulations to evaluate the proposed algorithm under various settings. Numerical results demonstrate that our algorithm achieves significant improvement over benchmark algorithms.
@@ II Related Work
User association is an effective approach to balance the workload among multiple BSs, hence improve the overall system performance. In [5], the authors formulate the user association problem as an integer programming prolem and develop an efficient iterative algorithm by relaxing the integral constraints. However, the obtained solution allows users to associate with multiple BSs, which is usually not realistic in practice. The authors in [6] consider the user association problem in the context of massive multiple-input-multiple-output networks. The objective is to maximize the network utility by optimizing the fraction of time that a user associate with each BS. This problem is further extended to the case with cooperative BSs in [7]. Although user association is well studied in the conventional communication paradigm [8, 9, 10], these methods are not applicable in semantic communication because we also need to consider the influence of users’ computational capabilities on their transmission demands.
The concept of semantic communication has been proposed for a long time [11]. Recently, with the advancement of artificial intelligence technology, there has been a growing interest in utilizing DNNs to implement semantic communication. The authors in [12] designed a semantic communication framework for text transmission. Their approach is extended to construct semantic communication systems for other types of sources (e.g. image[13], speech[14], and video[15]), which has shown significant improvements over the conventional communication paradigm. To improve the performance of semantic communication systems, the authors in [16, 17, 18, 19, 20, 21] investigate the optimal resource allocation in various situations. However, they did not take into account the influence of users’ computational capabilities. Besides, most of these works only consider a single BS, which may lead to suboptimal result in ultra-dense networks.
@@ III System Model and Problem Formulation
In this section, we first present the considered system and describe the concept of adaptive semantic communication. After that, we formulate an utility maximization problem and prove it is strongly NP-hard.
@@ III-A Network Model
As shown in Fig. 1, we consider a multi-cell network that consists of $M$ BSs and $N$ wireless devices (WDs), where one WD may communicate with multiple BSs. The set of all BSs and WDs are denoted by $\mathcal{M}$ and $\mathcal{N}$, respectively. There are a set $\mathcal{A}$ of $A$ applications running in the system. Let $\mathcal{N}_{a}\subseteq\mathcal{N}$ be the set of WDs belonging to application $a\in\mathcal{A}$. Each WD in $\mathcal{N}_{a}$ has a mount of raw data to be transmitted to BSs for further processing. For example, the WDs may be surveillance cameras that send captured photos to BSs for security risk detection. Due to limited wireless resources, the WDs will first extract the small-sized semantic information from the raw data (e.g. feature maps of the original photos) and only send the extracted semantic information to BSs.
In our considered model, a WD may be within the coverage of multiple BSs, so we need to determine which BS it communicates with. Specifically, the user association is denoted by the binary variable $x^{m}_{n}$, where $x^{m}_{n}=1$ if WD $n$ is associated with BS $m$, and $x^{m}_{n}=0$ otherwise. We assume a WD can communicate with at most one BS so $x^{m}_{n}$ should satisfy
@@ III-B Adaptive Semantic Communication Model
In existing semantic communication works, it is commonly assumed that the same transmitter is shared by all WDs. However, due to user heterogeneity, different WDs may prefer different semantic communication schemes. For example, if a WD has weak computational power but good channel conditions and rich spectrum resources, it may transmit the raw data to the BS without extracting the semantic information. In contrast, if a WD has strong computational power but limited communication resources, it can fully process the raw data and only send the final computation results to the BS. Therefore, by choosing appropriate semantic communication schemes, we can optimize the utilization of communication and computational resources, thereby enhancing system performance.
In practice, there are several ways to generate different semantic communication schemes. For example, the works in [22, 23, 24, 25] proposed various techniques to train DNNs of different model sizes in a single training process. Smaller models require less computation but also have worse accuracy. Since the semantic transmitters are implemented by DNNs, we can use these techniques to obtain a series of transmitters to achieve efficiency-accuracy trade-offs. Another method to derive multiple semantic communication schemes is through semantic compression [18, 19, 20, 21]. The basic idea is, after extracting the semantic information, the transmitter evaluates the importance of each component of the extracted semantic information and only transmit those that contribute most to the result accuracy. The evaluation of semantic information induces extra computation but the amount of data to be transmitted can be significantly reduced.
@@
According to the above discussion, for a given application $a$, its result accuracy is determined by the computation workload $c^{a}$ (defined as the number of CPU cycles required to extract the semantic information) and the communication workload $d^{a}$ (defined as the amount of data to be transmitted). Therefore, we can use the utility function $u^{a}(c,d)$ to represent the obtained utility when the computation and communication workload is $c$ and $d$. It is worth noting that in practice, $c$ and $d$ are often limited to discrete values. For example, in semantic compression, each component of the semantic information cannot be further divided so $d$ must be an integer multiple of the data size of each component. However, the possible values of $c$ and $d$ are usually relatively dense so $u^{a}(c,d)$ can be regarded as a sufficiently good approximation.
III-C Computation and Transmission Model
@@
In current practice, semantic information extraction is usually performed by DNNs, which requires a large amount of computation. For energy efficiency considerations, we assume WDs adopt the Dynamic Voltage and Frequency Scaling (DVFS) technique [26] to dynamically adjust the CPU frequency. For WD $n\in\mathcal{N}_{a}$, let $f_{n}$ be its CPU frequency and $c_{n}$ be its computation workload, then the computation time for semantic information extraction is
@@
According to the results in [27], the energy consumption per CPU cycle is proportional to the square of the frequency, which can be expressed as $\gamma_{n}f^{2}_{n}$, where $\gamma_{n}$ is the energy-efficiency factor determined by the chip architecture. Hence, the energy consumption due to computation is
@@
The spectrum bandwidth of each BS is equally divided into resource blocks (RBs) with bandwidth $W$ and the number of available RBs at BS $m$ is $K_{m}$. Let $z^{m}_{n}\in\{0,1,\dots,K_{m}\}$ be the number of BS $m$’s RBs allocated to WD $n$. As in [28, 29, 30, 31], we assume the inter-BS interference is randomized and/or mitigated by various coordination techniques so the total interference seen by each BS can be regarded as static. According to the Shannon capacity formula, the transmission rate from WD $n$ to BS $m$ can be expressed as:
@@
where $P_{n}$ is the transmission power of WD $n$, $h^{m}_{n}$ is the channel gain between WD $n$ and BS $m$, $\sigma^{2}$ is the noise power at BS $m$, and $I_{m}$ is the average interference seen by BS $m$. Clearly, for each BS $m$, the number of totally allocated RBs cannot exceed $K_{m}$, so we have
@@
For WD $n\in\mathcal{N}_{a}$, let $d_{n}$ be the amount of data to be transmitted, then the transmission time is
@@
and the energy consumption due to transmission is

### 09 Mar 2026 v3

# Joint User Association and Resource Allocation for Adaptive Semantic Communication in 5G and Beyond Networks

@@ Abstract
Semantic communication (SemCom) has emerged as a promising paradigm that leverages Deep Neural Networks (DNNs) to extract task-relevant information, thereby substantially reducing the volume of transmitted data. In existing implementations, the semantic transceiver is typically pre-trained for a specific task and uniformly adopted by all users. However, due to user heterogeneity in computational and communication capabilities, employing a single, fixed semantic transceiver may degrade the coding efficiency and transmission robustness. To address this issue, we first demonstrate the feasibility of dynamically adjusting the computational and communication overhead of DNN-based semantic transceivers, enabling a more flexible paradigm referred to as Adaptive Semantic Communication (ASC). Building on this concept, we formulate a joint user association and resource allocation problem for ASC in 5G and beyond networks, aiming to maximize overall system utility under energy and latency constraints. However, the problem is very challenging due to the inherent interdependencies among decision variables. To tackle this complexity, we decompose the original problem into three subproblems: (i) ASC scheme selection for each user, (ii) spectrum allocation at each Small-cell Base Station (SBS), and (iii) user association across SBSs. Each subproblem is solved sequentially based on the solutions of the preceding stages. The proposed algorithm efficiently yields near-optimal solutions with polynomial-time complexity. Simulation results demonstrate our approach outperforms existing baselines under various situations.
@@ I Introduction
The rapid proliferation of data-intensive applications—such as virtual reality, autonomous systems, and smart cities—has resulted in an unprecedented surge in data demand [2, 1]. Owing to the intrinsic limitations of communication resources (e.g., spectrum and time), such demand has already surpassed the transmission capacity of the conventional communication paradigm, which primarily emphasizes the accurate delivery of raw bits. To address this challenge, recent research has investigated innovative approaches that reduce the transmission volume through the integration of Artificial Intelligence (AI) techniques. In particular, a novel communication paradigm known as semantic communication (SemCom) has been proposed [4, 26, 12]. By leveraging AI, SemCom focuses on transmitting only the information that is meaningful and relevant to downstream tasks—commonly referred to as semantic information—thereby enabling substantial reductions in data traffic without compromising task performance.
Although SemCom effectively reduces the volume of transmitted data, it also introduces non-negligible computational overhead due to the extraction of semantic information. In current implementations, users performing the same downstream task typically share a common SemCom transceiver. However, due to the inherent heterogeneity in user devices and network conditions, it is desirable to dynamically adapt the transceiver based on the available computational and communication resources. For instance, as semantic transceivers are commonly realized using Deep Neural Networks (DNNs), we can actively reduce the DNN model size to mitigate computational demands for users with constrained computational capabilities. Similarly, for users experiencing poor channel conditions, the volume of transmitted data can be further decreased by selectively transmitting a subset of the semantic information. By jointly applying these two strategies, a spectrum of SemCom schemes can be constructed, each characterized by different levels of computational complexity, communication overhead, and inference accuracy. Selecting the most appropriate scheme for each user leads to the concept of Adaptive Semantic Communication (ASC), which significantly enhances the efficiency of both communication and computaional resources.
To demonstrate the potential of ASC in future communication systems, this paper investigates the performance optimization of ASC in 5G and beyond networks, where Small-cell Base Stations (SBSs) are densely deployed to meet the high-quality transmission demands of users. The proposed optimization framework simultaneously addresses user association, resource allocation, and ASC scheme selection to maximize system utility under energy and latency constraints. However, this problem is inherently challenging due to the complex interdependencies among decision variables. The primary challenges include:
  • Computation-communication trade-off: For each user, both extraction and transmission of semantic information consume time and energy. With limited time and energy budgets, a fundamental trade-off exists between the computational and communication workloads that a user can handle. To identify the optimal ASC scheme, it is crucial to characterize this trade-off and understand its effect on the inference accuracy of downstream tasks.
  • Spectrum sharing among heterogeneous users: Users associated with the same SBS must share limited spectral resources. However, due to user heterogeneity—such as variations in channel conditions, utility functions, and computational capabilities—the marginal utility of spectrum differs across users, making optimal spectrum allocation a non-trivial problem.
  • User association under spatial and network heterogeneity: In 5G and beyond networks, individual users are typically within the coverage of multiple SBSs. However, associating each user with its nearest SBS can lead to significant load imbalance due to the uneven spatial distribution of users and SBSs. Hence, optimizing user association requires a joint consideration of interrelated factors such as load balancing, channel conditions, and user-specific transmission demands.
  • Coupled decision variables across network layers: The decisions variables at different network levels are also deeply coupled. For example, user association determines the grouping of users that share spectrum resources. Furthermore, the spectrum allocation impacts each user’s transmission capacity, which influences the trade-off between communication and computation, and ultimately affects the choice of ASC schemes.
To address the aforementioned challenges, we decompose the original optimization problem into three hierarchical subproblems, each corresponding to a different network layer:
  • ASC Scheme Selection: We first focus on selecting the optimal ASC scheme for each user under fixed user association and spectrum allocation. To characterize the computation-communication trade-off, we formally prove that both energy and latency constraints are tight at the optimal solution. This key property enables us to efficiently determine the maximum transmittable data given a fixed computational workload. Building upon this result, we can derive the optimal computational workload and subsequently determine the optimal ASC scheme.
  • Spectrum Allocation: Next, we address the spectrum allocation under fixed user association. We consider two classes of utility functions: concave and general. For the concave case, we demonstrate that a simple greedy strategy yields the optimal allocation. For general utility functions, we develop a dynamic programming-based approach to efficiently determine the optimal solution.
  • User Association: Finally, we optimize user association across SBSs. The proposed method iteratively updates each user’s association starting from a dummy initialization. To mitigate computational complexity, we employ a linear approximation to estimate the marginal utility change for each potential reassignment. Furthermore, we introduce a kurtosis-based heuristic to determine the processing order of users.
These subproblems are inherently interrelated, with each depending on the optimal solutions of the preceding layers. By systematically leveraging the structural characteristics of each subproblem and their interconnections, we eventually obtain an efficient algorithm capable of producing near-optimal solutions with polynomial time complexity. The main contributions of this paper are summarized as follows:
  • We introduce the concept of ASC and discuss how to implement it in practice. Building on this foundation, we formulate a performance optimization problem for ASC in 5G and beyond networks, aiming to maximize the aggregate user utility subject to energy and delay constraints. Through formal analysis, we prove that the formulated problem is strongly NP-hard, highlighting its computational intractability.
  • To address the complexity of the original problem, we decompose it into three interrelated subproblems: ASC scheme selection, spectrum allocation, and user association. By identifying the mutual dependencies among the optimal solutions of these subproblems, we develop an efficient algorithm capable of achieving near-optimal solutions within polynomial time complexity.
  • We conduct extensive simulations to compare the proposed algorithm with both the conventional communication paradigm and existing SemCom schemes. The numerical results demonstrate that our approach significantly outperforms these benchmarks under diverse system configurations.
@@ II Related Work
II-A User Association in 5G and Beyond Networks
To meet the continuous growth in data traffic, telecommunication operators have been compelled to deploy a large number of SBSs, giving rise to the concept of Ultra-Dense Networks (UDNs) and Heterogeneous Netoworks (HetNets) [16, 30, 9, 36]. Given the limited communication resources available at each SBS, the design of effective user association strategies is crucial for balancing network load across SBSs and enhancing overall system performance [3, 21].
In [5], the authors formulated a user association problem aimed at maximizing network utility by optimizing the fraction of time each user is associated with different SBSs. This work was later extended to cooperative SBS scenarios in [40]. To address dynamic network conditions, a robust user association algorithm was proposed in [19] based on predicted future traffic fluctuations. The long-term user association problem was studied in [44], where a matching game-based algorithm was developed to efficiently maximize data rate and minimize the number of handovers.
User association is also commonly studied in conjunction with resource allocation—such as power and bandwidth—to achieve improved system performance [45, 25, 13, 29]. More recently, AI techniques have been introduced to solve the user association problem. Specifically, [24] proposed a decentralized multi-agent actor-critic framework to maximize energy efficiency in UDNs, while [37] applied a deep Q-network to learn optimal user association strategies from historical experience. In [18], a deep learning-based approach was developed to jointly optimize transmit power and user association.
Although user association is extensively investigated in existing literature, these studies are largely confined to the conventional communication paradigm. In contrast, our work introduces the ASC framework, which requires joint consideration of communication and computational resources, as well as their impact on the inference accuracy of downstream tasks. This added complexity significantly differentiates our problem from traditional user association formulations, thereby necessitating the development of novel algorithmic approaches.
II-B Resource Allocation for Semantic Communication
The concept of SemCom was originally introduced decades ago [28], but practical implementation remained elusive until recent advances in AI technologies enabled new possibilities. In particular, the emergence of DNNs has spurred growing interest in leveraging AI to realize SemCom systems. In [35], the authors proposed a pioneering SemCom framework for text transmission, which was subsequently extended to other modalities, including images [14], speech [33], and video [15]. These implementations demonstrated substantial performance gains compared to conventional communication paradigms.
To further enhance SemCom systems, several studies have explored semantic-aware resource allocation. For example, [39] investigated spectral efficiency in the semantic domain and re-examined resource allocation from a semantic perspective. In [38], the authors introduced the concept of semantic entropy to quantify task-relevant semantic information. They also developed a novel Quality-of-Experience (QoE) model for semantic-aware resource optimization and proposd a deep Q-network-based solution. Similarly, [47] designed an improved deep reinforcement learning (DRL) algorithm to enable server coordination during training, thereby accelerating convergence to a near-global optimum. The authors in [34] develop semantic channel models with a new performance metric and systematically optimize user association and bandwidth allocation. The work in [10] introduces a SemCom–empowered AIGC framework that leverages a resource-aware workload trade-off scheme to improve latency and content quality in dynamic wireless environments.
Recent research has also recognized that not all semantic information contributes equally to the inference accuracy of downstream tasks. As a result, prioritizing the transmission of high-impact semantic components has emerged as a promising strategy. For instance, [31] proposed a DRL-based algorithm using proximal policy optimization to jointly determine both resource allocation and the subset of semantic information to be transmitted. In [46], a dynamic resource allocation policy was designed using DRL to ensure that data with higher semantic value receives preferential access to limited communication resources. The study in [20] further explored the joint optimization of semantic information compression and resource block allocation to maximize task success probability.
The aforementioned studies primarily focus on the communication process, aiming to adaptively adjust communication overhead based on users’ available communication resources. However, in practical systems, users also exhibit significant heterogeneity in computational capabilities. The ASC framework proposed in this paper addresses this issue by simultaneously accommodating variations in both computational and communication resources across users. Building on this framework, we investigate the performance optimization of ASC in 5G and beyond networks, which necessitates the joint optimization of semantic scheme selection, user association, and resource allocation.
@@ III System Model and Problem Formulation
In this section, we first present the considered system and describe the concept of ASC. After that, we formulate an utility maximization problem and prove it is strongly NP-hard. The major notations used in this paper are summarized in Table I.
TABLE I: Major Notations
Notation Description
$M$ Number of SBSs.
NN Number of WDs.
AA Number of applications.
$x^{m}_{n}$ Whether WD $n$ is associated with SBS $m$.
$z^{m}_{n}$ Number of SBS $m$’s RBs allocated to WD $n$.
$c_{n},d_{n}$ Computation and communication workload of WD $n$.
$f_{n}$ CPU frequency of WD $n$.
$P_{n}$ Transmission power of WD $n$.
$u^{a}(c,d)$ Utility function of application $a$.
$h^{m}_{n}$ Channel gain between WD $n$ and SBS $m$.
$W$ Spectrum bandwidth of each RB.
$K_{m}$ Number of RBs available at SBS $m$.
$I_{m}$ Average interference at SBS $m$.
$r^{m}_{n}$ Transmission rate from WD $n$ to SBS $m$.
$T^{c}_{n},T^{t}_{n}$ Time consumption due to computation and transmission.
$E^{c}_{n},E^{t}_{n}$ Energy consumption due to computation and transmission.
@@ III-A Network Model
As illustrated in Fig. 1, we consider a 5G and beyond network comprising $M$ SBSs and $N$ wireless devices (WDs), where each WD may reside within the coverage area of multiple SBSs. The sets of SBSs and WDs are denoted by $\mathcal{M}$ and $\mathcal{N}$, respectively. The WDs may belong to various applications [^1] In this paper, application and downstream task are used interchangeably., such as intelligent surveillance and smart transportation. Let $\mathcal{A}$ represent the set of $A$ applications deployed in the system, and $\mathcal{N}_{a}\subseteq\mathcal{N}$ denote the subset of WDs associated with application $a\in\mathcal{A}$.
Each WD in $\mathcal{N}_{a}$ generates raw data that must be transmitted to the SBSs for further processing. For instance, WDs may consist of surveillance cameras that upload captured images for security threat detection. Given the constraints of limited wireless resources, WDs perform semantic extraction locally to reduce data volume. Specifically, instead of transmitting raw data, WDs extract compact semantic representations (e.g., feature maps of the original images) and transmit only this semantic information to the SBSs.
In our model, each WD may reside within the coverage area of multiple SBSs, so we need to decide which SBS the WD should associate with for communication. The user association is represented by a binary variable $x^{m}_{n}$, where $x^{m}_{n}=1$ indicates that WD $n$ is associated with SBS $m$, and $x^{m}_{n}=0$ otherwise. We assume that each WD can be associated with at most one SBS. Accordingly, the user association variables must satisfy the following constraint:
@@ III-B Adaptive Semantic Communication Model
In SemCom systems, the desired semantic information varies across different downstream tasks. As a result, a DNN-based semantic transceiver has to be trained for each specific downstream task. Once trained, the transceiver is deployed to all WDs performing the corresponding task. However, due to the inherent heterogeneity in user capabilities, different WDs may prefer transceivers with varying computational and communication demands.
Consider two illustrative extreme cases: a WD with limited computational capacity but favorable channel conditions and abundant spectrum resources may find it preferable to transmit raw data directly to the SBS, bypassing local semantic extraction. In contrast, a WD with strong computational capability but constrained communication resources may choose to fully process the data locally and transmit only the final inference result. By selecting appropriate SemCom schemes tailored to device capabilities, the system can achieve more efficient utilization of both communication and computational resources, thereby improving overall performance.
In practical implementations, various techniques can be employed to generate multiple SemCom schemes. For instance, methods proposed in [43, 41, 42, 7] enable training of DNNs with variable model sizes within a single training process. Smaller models reduce computational cost at the expense of lower inference accuracy. Given that semantic transceivers are implemented using DNNs, such techniques can be directly applied to produce a series of transceivers offering different trade-offs between efficiency and accuracy.
An alternative approach to constructing diverse SemCom schemes involves semantic compression [31, 46, 47, 20]. This technique involves evaluating the significance of individual components within the extracted semantic information and selectively transmitting only those that contribute most to the accuracy of downstream inference. Although semantic compression introduces additional computational overhead for evaluating component importance, it can substantially reduce the amount of data to be transmitted.
@@
Based on the above discussion, for a given application $a$, the resulting inference accuracy depends on the computation workload $c$ (defined as the number of CPU cycles required for semantic extraction) and the communication workload $d$ (defined as the volume of data to be transmitted). Accordingly, we define a utility function $u^{a}(c,d)$ to represent the utility derived from using a SemCom scheme with workloads $c$ and $d$. It is important to note that in practice, the values of $c$ and $d$ are typically discrete. For example, in semantic compression, the semantic information is composed of discrete components that cannot be further subdivided, implying that $d$ must be an integer multiple of the size of each component. Nevertheless, since the feasible values of $c$ and $d$ are often densely distributed, the utility function $u^{a}(c,d)$ can still serve as a sufficiently accurate approximation in optimization formulations.
III-C Computation and Communication Model
@@
In current practice, semantic information extraction is typically performed using DNNs, which entail substantial computational workloads. To improve energy efficiency, we assume that WDs employ the dynamic voltage and frequency scaling technique [32], which enables dynamic adjustment of CPU frequency based on processing demands. For WD $n\in\mathcal{N}_{a}$, let $f_{n}$ be its CPU frequency and $c_{n}$ be its computational workload, measured in CPU cycles. Then the computation time required for semantic information extraction can be expressed as:
@@
As established in [8], the energy consumption per CPU cycle is proportional to the square of the frequency. Accordingly, the total energy consumption for computation is given by:
@@
where $\gamma_{n}$ is a hardware-dependent energy efficiency coefficient determined by the chip architecture.
On the communication side, we assume that the available spectrum at each SBS is equally divided into resource blocks (RBs) with a fixed bandwidth $W$. The total number of RBs available at SBS $m$ is denoted by $K_{m}$. Let $z^{m}_{n}\in\{0,1,\dots,K_{m}\}$ represent the number of SBS $m$’s RBs allocated to WD $n$. In addition, the uplink interference at the SBS is typically the superposition of a large number of signals from spatially distributed users. Due to high user density and effective interference mitigation techniques (e.g., SIC, ZF, MMSE), the impact of any single user’s transmit power on the total interference is often negligible. Therefore, it is reasonable to regard the uplink interference from other cells as static [17, 27, 23, 11]. Under this assumption, the achievable transmission rate from WD $n$ to SBS $m$ is given by the Shannon capacity formula:
@@
where $P_{n}$ is the transmission power of WD $n$, $h^{m}_{n}$ denotes the channel gain between WD $n$ and SBS $m$, $\sigma^{2}$ is the noise power at SBS $m$, and $I_{m}$ represents the average interference experienced by SBS $m$. To ensure resource feasibility, the total number of RBs allocated by each SBS must not exceed its capacity:
@@
For WD $n\in\mathcal{N}_{a}$, let $d_{n}$ be the amount of semantic data to be transmitted, then the transmission time can be calculated as:
@@
and the corresponding energy consumption is: